I need a correction to my answer
the problem is (4+6i)(3+2i)+4i - (1+i)/(3-2i)
my answer was (343i-1)/13

Answer by CharlesG2(834)

"I need a correction to my answer
the problem is (4+6i)(3+2i)+4i - (1+i)/(3-2i)
my answer was (343i-1)/13"

(4 + 6i)(3 + 2i) + 4i - (1 + i)/(3 - 2i)
12 + 8i + 18i + 12i^2 + 4i - [(1 + i)(3 + 2i)]/[(3 - 2i)(3 + 2i)]
(multiplied the 2 complex numbers with FOIL, First Outer Inner Last,
and multiplied top and bottom of the fraction
by the conjugate of its denominator)
12 + 30i - 12 - (3 + 2i + 3i + 2i^2)/(9 + 6i - 6i - 4i^2)
(simplified and also multiplied out the numerator and denominator of the fraction with FOIL, 12i^2 = -12)
30i - (3 - 2 + 5i)/(9 + 4)
(simplified, 2i^2 = -2, -4i^2 = 4)
30i - (1 + 5i)/13
(simplified)
(30i * 13 - 1 - 5i)/13
(multiplied 30i by 13/13)
(390i - 5i - 1)/13
(simplified)
(385i - 1)/13
in a + bi form this would be -1/13 + (385/13)i